Solve and show step by step solution.



1. The density of the solution is 0.982 g/mL and the density of water is 1.00



g/ml. If the solution is composed of 30.6 g NH3 in 81.3 g of H2O. Compute for;



a. mole fraction of NH3



b. molarity of NH3



c. molality of NH3



2. Given the reaction, H3PO4 + 3 NaOH Na3PO4 + 3 H2O, what is the



mass of 0.20 M H3PO4 needed to react with 100 ml. of 0.10 M NaOH?

1
Expert's answer
2022-04-16T04:09:25-0400

"M(NH_3)=14+3x1=17 g\/mol"

"n(NH_3)=30.6\/17=1.8 mol"

"M(H_2O)=2x1+16=18 g\/mol"

"n(H_2O)=81.3\/18=4.52"

a. "w(NH_3)=1.8\/(4.52+1.8)=0.28"

b."V(H_2O)=81.3x1=81.3 ml=0.081 l"

"V(NH_3)=30.6\/0.982=31.2 ml=0.0312 l"

"c(NH_3)=1.8\/(0.081+0.0312)=16.1 mol\/l"

c. "c(NH_3)=1.8\/(0.0306+0.0813)=16.1 mol\/kg"


2. "n(NaOH)=0.1x0.1=0.01 mol"

"n(H_3PO_4)=0.01\/3=0.0033 mol"

"V(H_3PO_4)=0.0033x0.2=0.00067 l"

If dencity of H3PO4 is 1 g/ml, then

0.00067l=0.67 ml

0.67x1=0.67g